Adv. Math. Econ. 14, 99-112 (2010)

A certain Limit of Iterated Conditional Tail Expectation

Shigeo KUSUOKA ^{1}

Graduate School of Mathematical Sciences, The University of Tokyo, Komaba 3-8-1, Meguro-ku, Tokyo 153-8914, Japan

Received: June 18, 2009 Revised: October 21, 2009 JEL classification: C65. G19 Mathematical Subject Classification (2000): 60B05 Abstract. We consider the continuous limit of iterated Conditional Tail Expectation for a Brownian filtration and show the existence of the limit for some special random variables. Key words: Risk measure, Value at risk, CTE

1research supported by the 21st century COE project, Graduate School of Mathematical Sciences, The University of Tokyo